OpenStudy (anonymous):

16x^2 + 25 y ^ 2 - 32x + 50 y + 31 = 0

OpenStudy (anonymous):

combine like terms....first and what do you get?

OpenStudy (anonymous):

i get up to there but when i completed the square i was confused

OpenStudy (anonymous):

\[16(x ^{2}-2x)+25(y ^{2}+2y)+31=0\] applying completing the square method:: \[16((x ^{2}-2x+1)-1)+25((y ^{2}+2y+1)-1)+31=0\] \[16(x-1)^{2}+25(y+1)^{2}-16-25+31=0\] \[16(x-1)^{2}+25(y+1)^{2}=10\]

OpenStudy (anonymous):

perfect thats where i got up to

OpenStudy (anonymous):

now divide by 10 on both sides

OpenStudy (anonymous):

ya/...

OpenStudy (anonymous):

thats a problem

OpenStudy (anonymous):

\[(8/5)(x-1)^{2}+(5/2)(y+1)^{2}=1\] or \[(x-1)^{2}/(5/8)+(y+1)^{2}/(5/2)=1\] which is an ellipse

OpenStudy (anonymous):

ok but what's gonna be a ?

OpenStudy (anonymous):

a=\[\sqrt{5/8}\]

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

idk